OpenStudy (anonymous):

Ok, the problem asks to graph the line passing through the given point and having the given slope. Give the slope-intercept form of the equation of the line if possible. I don't understand this at all, no matter how much I work on it. (1,-1), m=-2/3

OpenStudy (anonymous):

us the info to right the equation that you will use to draw the line y=mx+c so m= -2/3 then y=(-2/3)x+c then you want the c value so put the coordinate in to the equation (1, -1) y= -1 x=1 -1=(-2/3)x1+c -1+(2/3)=c -1/3=c so thut the values together and you get y=(-2/3)x-(1/3)

OpenStudy (anonymous):

Still does not make sense at all.

OpenStudy (anonymous):

now that you have this equation you can plot the graph

OpenStudy (anonymous):

It says it is wrong every time I plot it.

OpenStudy (anonymous):

what is saying that it is wrong?

OpenStudy (anonymous):

I have no clue what I am doing wrong on the plotting. I put one point at -2/3 and the other one at 1/3.

OpenStudy (anonymous):

no thats not what you should do. first you create a domain for your x value. this is a list of x values you want to use. then put each x value in the equation and solve for y.